### v1

@@ Abstract
footnotetext: 1Korea University. 2Arizona State University. 3Meta. Work done while at ASU. Correspondence to: Young Geun Kim <younggeun_kim@korea.ac.kr>.
@@ 1 Introduction
Figure 1: End-to-end ISP pipeline for a vision DNN. Devices generate heterogeneous data due to the SW and HW variations, degrading model performance significantly.The accuracy of Homogeneous Client is obtained when FL devices are the same type, whereas Heterogeneous Client is obtained with different devices types.
@@
Despite the significant impact, system-induced data heterogeneity has been overlooked in FL research. Most prior works have primarily focused on label distribution skew, where a non-IID dataset is formed by partitioning a flat existing dataset based on the labels Li et al. (2020); Zhao et al. (2018); Mansour et al. (2020); Li et al. (2019); McMahan et al. (2017); Karimireddy et al. (2020). Hence, they cannot consider the characterization of cross-client heterogeneity which should be inherently included in the real-world FL datasets while significantly degrading the model performance.
Furthermore, such system-induced data heterogeneity deteriorates two notable problems in FL: fairness and domain generalization. In FL, fairness involves ensuring that the global-model fairly represents the learning from all devices and is not biased towards any particular device or group of devices Pessach & Shmueli (2022); Maeng et al. (2022). Domain generalization (DG), meanwhile, involves designing machine learning models that can generalize well to new, unseen domains Wang et al. (2022). In the context of FL, the domain can be a particular type of devices, each displaying unique characteristics due to system-induced data heterogeneity.
In this paper, we first characterize the impact of system-induced data heterogeneity for FL. We collect a vision dataset using a collection of devices of heterogeneous hardware [^1] The dataset will be open-sourced upon the acceptance to encourage further research in this domain.. By using the collected dataset, we analyze the variations and inconsistencies in the global model that arise due to system-induced data heterogeneity originated from hardware (HW) (i.e., sensors) and software (i.e., ISP algorithms) heterogeneity. We also investigate fairness and domain generalization implications which can stem from system-induced data heterogeneity in realistic FL environments. Based on the characterization results, we propose HeteroSwitch — a selective generalization method that combine the ISP transformation and SWAD — which counteracts the effects of system-induced data heterogeneity in FL. Compared with the baseline, HeteroSwitch reduces the variance of accuracy across device types by 79.5%, while improving the worst out-of-distribution (OOD) accuracy by 5.8%.
@@ 3.4 Deeper Look at Heterogeneity: SW
ISP algorithms are applied to RAW data to produce human visible images Hansen et al. (2021). To quantify the contributions of the ISP algorithm variations on system-induced data heterogeneity, we divide the ISP process into six primary stages, from demosaicing to compression, creating distinct images at each stage Buckler et al. (2017). Table 3 lists the algorithms used at each stage. The impact of each stage on model accuracy is assessed by either omitting a specific stage or applying another algorithm for each stage [^2] In case of demosaicing, which is a prerequisite for subsequent stages, we use two different methods rather than omitting this stage. — we train the global model using data processed by the Baseline column in Table 3, and test the model while adopting either Option 1 or Option 2 for each stage.
@@ 4.1 Fairness
Fig. 4 shows the model quality degradation for each device compared to the dominant devices — the x-axis represents the device type the model is deployed. As shown in Fig. 4, the accuracy on the 7 less dominant devices is 3.2% to 16.9% lower than that on the dominant devices. This implies that the global model has a bias toward the HW specification and SW algorithms of the dominant devices. On the other hand, although Galaxy S22 is the third most used for training, deploying the model to it yields the lowest accuracy among the devices. This also implies that a higher participation rate does not always guarantee a high accuracy, meaning that there exist the system features, such as the advanced ISP algorithms, which can smooth out (or deteriorate) the bias impact.
@@ 4.2 Domain Generalization
Fig. 5 shows the model quality degradation of global model accuracy when the device is excluded from training, compared to when all device types equally participate in FL (this practical is widely used in DG Dou et al. (2019); Segu et al. (2023)). The x-axis represents the device type that was excluded from training, while the y-axis shows the model quality degradation on the excluded (i.e., unseen) device. In DG, it is typically expected that a domain excluded from training would show lower accuracy, since the global model lacks exposure to its specific characteristics Dou et al. (2019); Segu et al. (2023). Interestingly, excluding a device from training does not consistently affect accuracy. For example, the accuracy for S9 drops when it is not part of the training. Conversely, devices like S6 and VELVET exhibit increased accuracy, even though they were excluded from training. This inconsistent result demonstrates system-induced data heterogeneity can deteriorate the complexity of DG in FL.
@@ 5.2 Generalization Techniques
Dataset Diversification with ISP Adjustment: Given heterogeneity across the data from different device types, we propose to expand the diversity of the data via random transformation — such transformation, including geometric, color adjustments, and random erasing, does not change the inherent label or meaning of the image, encouraging the model to learn from more varied range of data Shorten & Khoshgoftaar (2019).
Motivated by our observations in Section 3, which highlighted the need for addressing Tone and Color (especially White Balance Wyszecki & Stiles (2000)) variations, we suggest introducing random White Balance and tone transformations (via random gamma application Wyszecki & Stiles (2000)) to the dataset of each client.
@@ 6 Evaluation Results and Analysis
We configure FL experiments with total devices $N=100$, minibatch size $B=10$, selected devices for training for each round $K=20$, local epoch $E=1$, and number of communication rounds $T=1000$. For a DNN model, we use MobileNetv3-small, which is widely used in mobile execution environments Howard et al. (2019); Qian et al. (2021). We implement HeteroSwitch using PyTorch, and compare it with a baseline method, FedAvg McMahan et al. (2017), and prior works for data heterogeneity: q-FedAvg Li et al. (2019), FedProx Li et al. (2020), and Scaffold Karimireddy et al. (2020).
@@ 6.2 Comparison with Prior Works
q-FedAvg aims to minimize the accuracy variance among clients owing to data heterogeneity, by assigning weights to each client dataset based on the loss. However, q-FedAvg does not consider system-induced data heterogeneity across different device types in FL. As a result, although q-FedAvg improves the variance and average accuracy by 37.4% and 1.5%, respectively, over the baseline, it fails to generalize to unseen device type with 2.9% decrease in worst-case accuracy.
FedProx mitigates data heterogeneity in FL by adjusting the magnitude of local updates with an additional L2 regularization. Similar to q-FedAvg, this method does not focus on system-induced data heterogeneity. Therefore, although FedProx improves the worst-case accuracy, variance and average accuracy by 0.5%, 2.3%, and 1.8%, respectively, over the baseline, it shows lower performance compared to HeteroSwitch.
@@ 7 Related Work
Data Heterogeneity in FL: Federated Learning, with an increased number of clients, inherently produces diverse data, reflecting variations in class distribution, participant geography, culture, and device usage patterns Kairouz et al. (2021), leading to data heterogeneity. Prior works have tried to mitigate this by adopting a regularization method to local models Li et al. (2020), sharing a small amount of public data across local clients Zhao et al. (2018); Mansour et al. (2020), or employing weighted averaging for model aggregation Li et al. (2019). To the best of our knowledge, our work is the first to analyze the problems caused by system-induced data heterogeneity in FL where a range of device types can participate in the learning process and propose a solution that addresses the aforementioned challenges while still preserving the principle of data privacy in FL.

### v2

@@ Abstract
footnotetext: 1Korea University. 2Arizona State University. 3Meta (work done while at ASU). Correspondence to: Young Geun Kim <younggeun_kim@korea.ac.kr>.
@@ 1 Introduction
Figure 1: End-to-end ISP pipeline for a vision DNN. Devices generate heterogeneous data due to the SW and HW variations, degrading model performance significantly.The accuracy of Homogeneous Client is obtained when FL devices are the same, whereas Heterogeneous Client is obtained with different device types.
@@
Despite the significant impact, system-induced data heterogeneity has been overlooked in FL research. Most prior works have primarily focused on label distribution skew, where a non-IID dataset is formed by partitioning a flat existing dataset based on the labels Li et al. (2020); Zhao et al. (2018); Mansour et al. (2020); Li et al. (2019); McMahan et al. (2017); Karimireddy et al. (2020). Hence, they cannot consider the characteristics of cross-client heterogeneity which should be inherently included in the real-world FL datasets significantly degrading the model performance.
Furthermore, such system-induced data heterogeneity deteriorates two notable problems in FL: fairness and domain generalization. In FL, fairness involves ensuring that the global-model fairly represents the learning from all devices and is not biased towards any particular group of devices Pessach & Shmueli (2022); Maeng et al. (2022). Domain generalization (DG), meanwhile, involves designing machine learning models that can generalize well to new, unseen domains Wang et al. (2022). In the context of FL, the domain can be a particular type of devices, each displaying unique characteristics due to system-induced data heterogeneity.
In this paper, we first characterize the impact of system-induced data heterogeneity for FL. We collect a vision dataset using a collection of devices of heterogeneous hardware [^1] The dataset is available at https://github.com/CASL-KU/HeteroSwitch.. By using the collected dataset, we analyze the variations and inconsistencies in the global model that arise due to system-induced data heterogeneity originated from hardware (i.e., sensors) and software (i.e., ISP algorithms). We also investigate fairness and domain generalization implications which can stem from system-induced data heterogeneity in realistic FL environments. Based on the characterization results, we propose HeteroSwitch — a selective generalization method that combines the ISP transformation and SWAD — which counteracts the effects of system-induced data heterogeneity in FL. Compared with the baseline, HeteroSwitch reduces the variance of accuracy across device types by 79.5%, while improving the worst out-of-distribution (OOD) accuracy by 5.8%.
@@ 3.4 Deeper Look at Heterogeneity: SW
ISP algorithms are applied to RAW data to produce human visible images Hansen et al. (2021). To quantify the contributions of the ISP algorithm variations on system-induced data heterogeneity, we divide the ISP process into six primary stages, from demosaicing to compression, creating distinct images at each stage Buckler et al. (2017). Table 3 lists the algorithms used at each stage. The impact of each stage on model accuracy is assessed by either omitting a specific stage or applying another algorithm for each stage [^2] In case of demosaicing, which is a prerequisite for subsequent stages, we use two different methods rather than omitting the stage. — we train the global model using data processed by the Baseline column in Table 3, and test the model while adopting either Option 1 or Option 2 for each stage.
@@ 4.1 Fairness
Fig. 4 shows the model quality degradation for each device compared to the dominant devices — the x-axis represents the device type the model is deployed. As shown in Fig. 4, the accuracy on the 7 less dominant devices is 3.2% to 16.9% lower than that on the dominant devices. This implies that the global model has a bias toward the HW specification and SW algorithms of the dominant devices. On the other hand, although Galaxy S22 is the third most used for training, deploying the model to it yields the lowest accuracy among the devices. This also implies that a higher participation rate does not always guarantee a high accuracy, meaning that there exist the system features, such as the advanced ISP algorithms, which can smooth out (or deteriorate) the bias.
@@ 4.2 Domain Generalization
Fig. 5 shows the model quality degradation of global model accuracy when the device is excluded from training, compared to when all device types equally participate in FL (this practical is widely used in DG Dou et al. (2019); Segu et al. (2023)). The x-axis represents the device type that was excluded from training, while the y-axis shows the model quality degradation on the excluded (i.e., unseen) device. In DG, it is typically expected that a domain excluded from training would show lower accuracy, since the global model lacks exposure to its specific characteristics Dou et al. (2019); Segu et al. (2023). Interestingly, excluding a device from training does not consistently affect accuracy as they generate different details for samples, such as color, texture, and contextual details Tommasi et al. (2017). For example, the accuracy for S9 drops when it is not part of the training.
Conversely, older devices like S6, Nexus5X and G4, which have lower resolutions and simpler ISP algorithms, commonly exhibit increased accuracy, even though they were excluded from training. This inconsistent result demonstrates system-induced data heterogeneity can deteriorate the complexity of DG in FL.
@@ 5.2 Generalization Techniques
Dataset Diversification with ISP Adjustment: Given heterogeneity across the data from different device types, we propose to expand the diversity of the data via random transformation during training — such transformation, including geometric, color adjustments, and random erasing, does not change the inherent label or meaning of the image, encouraging the model to learn from more varied range of data Shorten & Khoshgoftaar (2019).
Motivated by our observations in Section 3, which highlighted the need for addressing Tone and Color (especially White Balance Wyszecki & Stiles (2000)) variations, we temporarily adopt random White Balance and tone transformations (via random gamma application Wyszecki & Stiles (2000)) to the dataset of each client during local training.
@@ 6 Evaluation Results and Analysis
We configure FL experiments with total devices $N=100$, minibatch size $B=10$, selected devices for training for each round $K=20$, local epoch $E=1$, and number of rounds $T=1000$. For a DNN model, we use MobileNetv3-small Howard et al. (2019), which is widely used in mobile execution environments. We implement HeteroSwitch using PyTorch, and compare it with a baseline method, FedAvg McMahan et al. (2017), and prior works for data heterogeneity: q-FedAvg Li et al. (2019), FedProx Li et al. (2020), and Scaffold Karimireddy et al. (2020).
@@ 6.2 Comparison with Prior Works
q-FedAvg aims to minimize the accuracy variance among clients owing to data heterogeneity, by assigning weights to each client dataset based on the loss. However, q-FedAvg does not consider system-induced data heterogeneity across different device types in FL. As a result, it fails to generalize to unseen device type with 2.9% decrease in worst-case accuracy; it improves the variance and average accuracy by 37.4% and 1.5%, respectively, over the baseline though.
FedProx mitigates data heterogeneity in FL by adjusting the magnitude of local updates with an additional L2 regularization. Similar to q-FedAvg, this method does not focus on system-induced data heterogeneity. Hence, although it improves the worst-case accuracy, variance and average accuracy by 0.5%, 2.3%, and 1.8%, respectively, over the baseline, it shows lower performance compared to HeteroSwitch.
@@ 6.4 Impact on Realistic FL Dataset
6.5 Impact on Synthetic Dataset
In alignment with our findings on collected dataset, we further explore the impact of system-induced heterogeneity using the synthetic CIFAR-100 dataset Krizhevsky et al. (2009). To emulate the diverse characteristics observed in real device data (Section 3), we inject the heterogeneity into the CIFAR-100 dataset by implementing 10 different randomized settings for contrast, brightness, saturation, and hue. A simple CNN model is used to quantify the impact of these modifications in FL setting. The distribution of accuracy across synthetic device types with FedAvg and HeteroSwitch is illustrated in Figure 8.
Under synthetic condition, FedAvg achieves an average accuracy of 38.34% but exhibits a variance of 212.97 across the synthetic devices. This scenario underscores the challenges posed by system-induced data heterogeneity, as certain device types show decreased accuracy. In contrast, HeteroSwitch significantly outperforms FedAvg, enhancing accuracy by 24.4% and reducing variance by 43.9%, demonstrating its effectiveness in managing device heterogeneity.
Figure 8: Model accuracy on 10 different synthetically generated device types with CIFAR-100
6.6 Impact on Non-vision Dataset
We expand our evaluation to non-vision data, an Electrocardiogram (ECG) dataset. This dataset comprises data from four distinct sensor types, each introducing unique noise patterns and thereby contributing to heterogeneity in the dataset Vollmer et al. (2022). A simple DNN model, designed to calculate heart rates from ECG signals, is used to quantify the impact of sensor heterogeneity.
Using FedAvg, heart rate predictions from the same individual ECG data show a significant divergence, with an average deviation of 31.8% due to sensor variability. In contrast, HeteroSwitch, equipped with a Random-Gaussian Filter, notably reduces this deviation to 18.3%. This demonstrates that system-induced data heterogeneity is not confined to vision data and highlights the capability of HeteroSwitch to mitigate such heterogeneity across diverse data types.
@@ 7 Related Work
Data Heterogeneity in FL: Federated Learning, with an increased number of clients, inherently produces diverse data, reflecting variations in class distribution, participant geography, culture, and device usage patterns Kairouz et al. (2021), leading to data heterogeneity. Prior works have tried to mitigate this user-induced heterogeneity by adopting a regularization method to local models Li et al. (2020), sharing a small amount of public data across local clients Zhao et al. (2018); Mansour et al. (2020), or employing weighted averaging for model aggregation Li et al. (2019). Despite these efforts, differences between devices remain evident even after addressing the user-induced heterogeneity. Given that different causes of data heterogeneity result in unique characteristics in the feature space (where each potentially requires own distinct solution) we focus on the system-induced heterogeneity in this work. To the best of our knowledge, our work is the first to analyze the problems caused by system-induced data heterogeneity in FL where a range of device types can participate in the learning process and propose a solution that addresses the aforementioned challenges while still preserving the data privacy in FL.
@@ 8 Conclusion
Acknowledgements
This work was supported in part by National Research Foundation of Korea (NRF) grants funded by the Korea government (MSIT) (2021R1C1C1008617 and RS-2023-00212711), ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-2020-0-01819), and ITRC (Information Technology Research Center) support program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2024-RS-2023-00260091).